# A new hierarchy for complex plane curves

### v2

@@ 2.4. Exact sequences for line arrangements
$A$ a generating set for $M^{\prime}$ and a set $B\subset M$ which maps onto a generating set for $M^{\prime\prime}$, then the image of $A$ under $M^{\prime}\to M$ union $B$ gives a generating set for $M$. We apply this to $M^{\prime}=D(f^{\prime})$, $M=D(f)$ and $M^{\prime\prime}=D(f^{\prime\prime})$, and $A$ consisting of $E^{\prime}=E$ the Euler derivation of ${\mathcal{A}}^{\prime}$ and of ${\mathcal{A}}$, plus the two generators of $D(f^{\prime})_{0}$ and $B$ consisting of $E$ and an element ${\alpha}\in D(f)$ such that $\rho({\alpha})$ is the generator of $D(f^{\prime\prime})$ of degree $n-1$. The element $xE$ is discarded from the union since we have $E$ in this union.

### v3

@@ 2.4. Exact sequences for line arrangements
AA a generating set for $M^{\prime}$ and a set $B\subset M$ which maps onto a generating set for $M^{\prime\prime}$, then the image of $A$ under $M^{\prime}\to M$ union $B$ gives a generating set for $M$. We apply this to $M^{\prime}=D(f^{\prime})$, $M=D(f)$ and $M^{\prime\prime}=D(f^{\prime\prime})$, and $A$ consisting of $E^{\prime}=E$ the Euler derivation of $\mathcal{A}$$^{\prime}$ and of $\mathcal{A}$}, plus the two generators of $D(f^{\prime})_{0}$ and $B$ consisting of $E$ and an element ${\alpha}\in D(f)$ such that $\rho({\alpha})$ is the generator of $D(f^{\prime\prime})$ of degree n1n-1. The element $xE$ is discarded from the union since we have $E$ in this union.
@@ Acknowledgement
We would like to thank the referees for their careful reading of the manuscript and for their very useful suggestions to improve the presentation.
@@
Alexandru Dimca is partially supported by the project “Singularities and Applications” - CF 132/31.07.2023 funded by the European Union - NextGenerationEU - through Romania’s National Recovery and Resilience Plan.
@@
Conflict of Interests
We declare that there is no conflict of interest regarding the publication of this paper.
Data Availability Statement
We do not analyze or generate any datasets, because this work proceeds within a theoretical and mathematical approach.